All ISS systems continue to function nominally, except those noted previously or below. Sunday – Flight Day 10 (FD10) of STS-126/ULF-2. ISS crew work cycle today: Wake 8:55am EST; sleep 12:25am; Shuttle crew: 12:55am (both until 8:55am tomorrow). Ahead: Week 5 of Increment 18.
Mission ULF-2’s EVA-3 was completed successfully last night by Heidemarie Stefanyshyn-Piper & Steve Bowen in 6h 57min, accomplishing nearly all of its objectives and some get-aheads.
[During the spacewalk, Piper (EV1) & Bowen (EV2) –
Additionally, as get-ahead tasks for EVA-4, the spacewalkers removed covers 17 & 18 and TBA-3.
Cleaning and lubing of the race ring bearing surfaces beneath covers 17 & 18, plus replacement of TBA-3 will be completed during tomorrow’s EVA-4.
Official start time of the spacewalk was 1:01pm EST, and it ended at 7:58pm. Total EVA duration (PET = Phase Elapsed Time) was 6h 57min. It was the 117th spacewalk for ISS assembly & maintenance and the 89th from the station (65 from Quest, 24 from Pirs, plus 28 from Shuttle) totaling 551h, the third EVA for Expedition 18 and the 17th so far this year (including two Russian Orlan EVAs). It was the 5th spacewalk for Stefanyshyn-Piper (33h 42m tot.) and the second for Bowen (13h 49m tot.). After today's EVA, a total of 155 spacewalkers (119 NASA astronauts, 25 Russians, and 11 astronauts representing Japan-1, Canada-4, France-1, Germany-2 and Sweden-3) have logged 739h 22min outside the station on building, outfitting & servicing. It was the 137th spacewalk involving U.S. astronauts.]
After the spacewalkers’ ingress (7:58pm EST), post-EVA activities by MS3 Piper, MS2 Bowen, FE-2 Chamitoff, CDR Fincke & CDR Ferguson in the A/L and preparations for EVA-4 tomorrow consisted of --
In support of EVA-4, Sandra Magnus will set up the photo/TV infrared camera for TRJJ imagery.
CDR Fincke & FE-2-18 Magnus are continuing with WRS (Water Recovery System) activation activities. [The UPA (Urine Processor Assembly) was activated again but shut down after about 2 hrs due to low centrifuge speed, similar to yesterday’s shutdown. Engineering teams met and determined that a possible cause of the low centrifuge speed is a vibration-induced resonance. The on-orbit unit is mounted on vibration isolators. In a late-added task estimated at 2 hrs, Mike & Sandy are performing an on-orbit modification today to remove the vibration isolators and hard-mount the UPA on the rack structure.]
The WPA (Water Processor Assembly) completed its first on-orbit water processing, using about 40 lbs of waste water (mostly condensate). Magnus collected samples of this first WPA-produced water for archival storage & in-flight chemistry/microbiology analysis with the U.S. WMK (water microbiology kit) and one sample for the newly installed TOCA (Total Organic Carbon Analyzer), then downlinked the results of the latter via USB stick and SSC (Station Support Computer). [The first two replications of the analysis cycle were completed, but the analysis terminated during the third replication. The problem could have been caused by too much gas being pushed through the system which created bubbles. There is also a chance that an internal water leak is responsible. Another software update may be required.]
FE-1 Lonchakov conducted his first session with the Russian biomedical MBI-15 "Pilot-M"/NEURO signal response experiment after setting up the workplace and equipment. Afterwards, the Pilot-M & Neurolab-2000M gear was disassembled and stowed away. [MBI-15 requires a table, ankle restraint system, eyeball electrodes for an EOG (electrooculogram), and two hand controllers (RUO & RUD) for testing piloting skill in “flying” simulations on a laptop (RSK1) under stopwatch control, as well as for studying special features of the psychophysiologic response of cosmonauts to the effects of stress factors in flight.]
FE-2 Chamitoff set up the payload equipment for his fourth session with the 3D-Space (Mental Representation of Spatial Cues During Space Flight) experiment and then performed the protocol, with all three exercises (distance, illusion, handwriting). After all data were stored on PCMCIA memory card, the ESA multipurpose laptop was shut down and the hardware disconnected & stowed away. [3D-SPACE, a collaboration of ESA and the French space agency CNES (Centre National de la Recherche Scientifique), is designed to identify accurate visual perception & localization of objects in the space environment as prerequisites for spatial orientation & reliable performance of motor tasks in microgravity. Humans have mental representations of their surroundings based on sensory information & experience. It is hypothesized that depth & distance perception of objects could be altered in space due to the absence of gravitational reference and ambiguous perspective cues. 3D-SPACE investigates (a) depth perception & the role of perspective cues using geometric illusions, (b) distance perception using both standard psychophysics tests & natural three-dimensional scenes presented on a virtual reality head-mounted display, and (c) the effects of cognitive vs. perceptual-motor changes using handwriting & drawing tests.]
In the COL (Columbus Orbital Laboratory), Gregory installed the ITCS (Internal Thermal Control System) Sampling Adapter to retrieve a sample of its coolant for return to the ground in the Shuttle middeck.
Later today, the FE-2 will hook the AmiA (Antimicrobial Applicator) into the Kibo JPM (Japanese Pressurized Module)’s ITCS, connecting it to the MTL supply QD. [AmiA introduces OPA (Ortho-Phthalaldehyde), an antimicrobial agent, into the JPM TCS coolant. The applicator must remain installed in the loop for at least 6 hrs.]
Yesterday’s routing & installing of the PWD (Potable Water Dispenser) standoff hose by Greg & Sandy was not completed due other hoses obstructing the space required. [The hose is routed from the LAB1S7 rack location to the Aft Deck End Cone and from there to the Port Deck Standoff. The two FEs removed the obstruction by taking off a flange, then taped hose & flange to the front of the Z Panel on the rack. Specialists will determine the future plan. Temporary shutdown of the LTL (Low Temperature Loop) may help to get the hose installed.]
Sandra Magnus took the periodic O-OHA (on-orbit hearing assessment) test, a 30-min. NASA environmental health systems examination to assess the efficacy of acoustic countermeasures, using a special MEC (Medical Equipment Computer) laptop application. It was Sandy’s first O-OHA test. [The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, new Bose ANC headsets (delivered on 30P) and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC, featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed once per month. Note: There have been temporary hearing deficits documented on some U.S. and Russian crewmembers, all of which recovered to pre-mission levels.]
FE-2-18 Magnus performed outfitting in the Node -2 by installing a guide ring on the EVA side of the nadir hatch.
In response to a JAXA request, Magnus will also make a second attempt at retracting the structural latch (#1) of the EFBM (Exposed Facility Berthing Mechanism) by activating the BCDU (Berthing Mechanism Control & Display Unit) in the JPM. [During yesterday’s checkout of the Kibo EFBM by Mike & Sandy structural latch #1 failed to retract. The retraction was added as a 20-min contingency task to EVA-4.]
Preparatory to today’s scheduled waste water dump from the Shuttle, FE-2-18 Magnus verified closure of the protective window shutters in the Kibo JPM (Japanese Pressurized Module). Later today, CDR Ferguson will maneuver the ISS/Shuttle stack into the proper attitude at ~5:46pm EST, and at ~5:46pm the venting (in retrograde direction) will commence, lasting about 50 min. –XVV attitude (Shuttle bottom facing opposite to flight direction for TPS protection) will then be restored with Shuttle thrusters (ORB mode) at ~7:00pm and taken over by CMG Momentum Management. [This will be the last waste water dump for this mission, even if it is extended by a day.]
FE-1 Lonchakov performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM (Service Module), including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP-Moscow. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
SOZh checkups by Yuri today also include the Russian POTOK-150MK (150 micron) air filter unit of the SM’s SOGS air revitalization subsystem, gathering weekly data on total operating time & “On” durations for reporting to TsUP-Moscow.
Lonchakov is to conduct the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the passageways PrK (SM Transfer Compartment)–PrK–RO (SM Working Compartment), PkhO (SM Transfer Tunnel)–RO, PkhO–DC1, PkhO–FGB PGO, FGB PGO–FGB GA, FGB GA–Node-1. [This checkup is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).]
The FE-1 will also shoot situational awareness video of ventilation air ducts in the SM, DC1 Docking Compartment and Soyuz spacecraft, followed by downlinking the footage to TsUP-Moscow.
The ISS crew completed their physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2, FE-2-18), TVIS treadmill (CDR, FE-1), RED (CDR, FE-2), and VELO bike with bungee cord load trainer (FE-1). It was the second exercise session for Sandy Magnus, now to become a daily protocol for her.
At ~5:05pm EST, the combined crew is scheduled for three PAO TV downlinks with ABC News (Gina Sunseri), CBS News (Bill Harwood, Peter King) and NBC News (Lester Holt).
At ~8:50pm, the ten crewmembers are scheduled for an in-depth one-hour review of procedures for tomorrow’s 6:30h EVA-4 spacewalk, with egress tomorrow afternoon at ~1:45pm EST.
Bowen (EV2) and Kimbrough (EV3) will begin their “campout” in the “Quest” Airlock (A/L) tonight at ~11:20pm with hatch closure and depressurization of the CL (Crewlock) from 14.7 to 10.2 psi, followed by mask prebreathe until ~12:25m. Sleep time for the ISS crew begins at 12:25am, for the Shuttle crew half an hour later.
EVA-4 Look-ahead: Following the usual hygiene break/with mask prebreathe for Bowen & Kimbrough at ~9:30am-10:40am tomorrow morning after spending the night on 10.2 psi, the A/L hatch will be closed again for EVA preps in 10.2 psi, followed by EMU purge and prebreathe in the EMUs. Afterwards, with CL depressurization and EV1/EV2 egress, EVA-4 nominally begins at ~1:45pm EST. Objectives of the spacewalk are –
Possible get-aheads: S1 Radiator, P1 Radiator.
The NOAX (Non-Oxide Adhesive Experimental) caulk applicator from the Shuttle, using Braycote-601 lubricant for the trundle bearings, will be left in the A/L for use during EVA-4 if needed. It will not be modified to reduce Braycote flow from its nozzle
KURS-P Repair Update: After the K2-VKA-01 container of the KURS-P automated rendezvous radar system was installed and connected yesterday in the SM by FE-1 Lonchakov, it was successfully tested by TsUP-Moscow and is running with no issues. [The passive KURS-P is required for the Progress 31P docking next week (11/30) at the DC1 nadir port, with KURS backup provided by the manual TORU system.]
Cargo Transfers: MPLM (Multi-Purpose Logistics Module) cargo transfers by Pettit, Boe, Piper, Chamitoff, Magnus and Fincke are continuing at a brisk pace, remaining slightly ahead of schedule. [As of this morning, Middeck transfers were 56% complete, MPLM 84%, overall 76%.]
No CEO photo targets uplinked for today.
CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).
ISS Orbit (as of this morning, 11:55am EST [= epoch]):
Mean altitude -- 353.8 km
Apogee height -- 358.3 km
Perigee height -- 349.3 km
Period -- 91.62 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0006641
Solar Beta Angle -- 15.8 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.72
Mean altitude loss in the last 24 hours – 0 m
Revolutions since FGB/Zarya launch (Nov. 98) – 57363.
Significant Events Ahead (all dates Eastern Time, some changes possible.):
11/26/08 -- Progress M-66/31P launch
11/27/08 -- STS-126/Endeavour/ULF-2 undocking; 10:40am – Under Review
11/29/08 -- STS-126/Endeavour/ULF-2 landing; ~2:10pm – Under Review
11/30/08 -- Progress M-66/31P docking – DC1 Nadir (~7:23am)
12/07/08 -- Progress M-65/30P reentry (after 3 weeks autonomous flight for geophysical experiments)
12/18/08 -- Russian EVA-21
02/09/09 -- Progress M-66/31P undocking & deorbit
02/10/09 -- Progress M-67/32P launch
02/12/09 -- Progress M-67/32P docking
02/12/09 -- STS-119/Endeavour/15A launch – S6 truss segment
02/14/09 -- STS-119/Endeavour/15A docking
02/24/09 -- STS-119/Endeavour/15A undocking
02/26/09 -- STS-119/Endeavour/15A landing (nominal)
03/25/09 -- Soyuz TMA-14/18S launch
03/27/09 -- Soyuz TMA-14/18S docking (DC1)
04/05/09 -- Soyuz TMA-13/17S undocking
04/07/09 -- Progress M-67/32P undocking & deorbit
05/15/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD
05/27/09 -- Soyuz TMA-15/19S launch
Six-person crew on ISS
07/30/09 -- STS-128/Atlantis/17A – MPLM (P), last crew rotation
10/15/09 -- STS-129/Endeavour/ULF3 - ELC1, ELC2
12/10/09 -- STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 -- STS-131/Atlantis/19A – MPLM(P)
04/08/10 -- STS-132/Endeavour/ULF4 – ICC-VLD, MRM1 (contingency)
05/31/10 -- STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).